OpenStudy (anonymous):

The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units? No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64 No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32

OpenStudy (anonymous):

@ParthKohli @phi @ganeshie8

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

I have no idea how to math... @KingGeorge

OpenStudy (kinggeorge):

Remember that if you want to find the perimeter of a rectangle with length \(x\) and height \(y\), the formula is just\[P=2(x+y).\]In your case, you're already given that the perimeter \(P\) is \(64\). So if your values of \(x\) and \(y\) were possible, then\[64=2(x+y)\implies32=x+y\]So now, which of the options do you think you should choose?

OpenStudy (anonymous):

D?

OpenStudy (kinggeorge):

We need \(x+y\) to equal 32, not be less than 32.

OpenStudy (anonymous):

Oh okay, sorry I am a bit dumb, lol

OpenStudy (anonymous):

im gonna say b

OpenStudy (kinggeorge):

That's right. Excellent!

OpenStudy (anonymous):

Thanks so much!! :)